Is the following differential equation exact or not and solve it if its exact ?


( x^2 + cosy ) + ( y^2+x+sinx) dy/dx = x
1
Expert's answer
2012-03-15T10:11:58-0400
Differential equation of the form I(x,y)dx + J(x,y)dy = 0 is exact only if dI/dy =dJ/dx .
We can transform our equation to this form :
( x^2 + cosy )dx + ( y^2+x+sinx )dy = xdx
( x^2 - x + cosy )dx + ( y^2+x+sinx )dy = 0 .
In this equation I(x,y) = x^2 - x + cosy and J(x,y) = y^2+x+sinx .
Then we find : dI/dy = -siny and dJ/dx = 1 + cosx.
We see that dI/dy not equal dJ/dx .
That's why given equation is not exact.